### v1

@@ 3.4 Open Education
We are compiling a dynamic textbook titled “High-Performance Seismology" that the workshop participants have contributed to HPS.
@@ 4.1 Virtual Events
In April 2022, we organized a two-day workshop on moment tensor estimation using the open-source MTUQ software. The first day featured a 2-hour session introducing key concepts and tutorials. The second day consisted of a 4-hour session that demonstrated how to calculate a library of Green’s functions for a specified 1D layered model using a frequency-wavenumber code (Zhu and Rivera, 2002) and obtain a seismic moment tensor solution. Attendance was strong, with 78 on day 1 and 68 on day 2, indicating sustained interest in the more detailed content. In preparation, software containers for four systems (Windows/PC, Linux, Mac OS Intel and Mac OS Apple Silicon processors) were developed and tested, resulting in high success rates for participants running the examples.
@@
Building on the success of the previous workshop, we held a three-day SPECFEM users’ workshop in October 2022. Each of the three daily, 4-hour sessions had a specific focus: the forward wavefield (day 1), sensitivity kernels (day 2), and seismic imaging (day 3). Each session included short (20-minute) science lectures, 45-minute tutorials that participants could run locally using pre-downloaded software containers, and wrap-up discussion sessions. Participation ranged from 187 attendees in the day 1 opening seminar to 63 in the day 3 discussion (Figure 5). This was the first SPECFEM workshop featuring seismic imaging, providing a natural progression from synthetic seismograms (day 1) to sensitivity kernels (day 2) to iterative tomographic inversion using SeisFlows and Pyatoa (day 3) (Modrak et al., 2018; Chow et al., 2020). Crafting a pedagogical but research-grade notebook took dedicated efforts.
@@
In 2023, we held a four-day virtual training workshop in collaboration with the SCEC and several European and NSF-funded projects. Each day focused on a specific theme, starting with an opening day of lectures on open science, reproducibility, software best practices, and an introduction to HPC and Cloud Computing. Subsequent days were divided into subdisciplines and platforms. The workshop attracted over 200 interested participants, with 80 joining on Zoom at the workshop’s start, though attendance varied due to time zone challenges. Designed for tool adoptions and relatively fast-paced, participants were exposed to diverse topics in seismology, including earthquake simulations focusing on dynamic rupture (with SeisSol, Käser et al. (2010); Uphoff et al. (2024)) and wave propagation (SPECFEM, Komatitsch et al. (2002)), machine learning phase picking (ELEP, Yuan et al. (2023)), earthquake probabilistic forecasting (pyCSEP, Savran et al. (2022)), and user access to Community Earth Models maintained by SCEC, (e.g., Plesch et al., 2007; Small et al., 2017).
@@
The 2024 MsPASS training short course was hosted in collaboration with EarthScope during the week of July 8, 2024, as part of their 2024 Technical Short Course series. The event featured two hours of lectures and hands-on sessions over three days. Participants could attend the course in real time or access recordings on YouTube afterward. Daily homework assignments were given, and an optional final project. The application-based enrollment process received 99 valid applications, from which 53 participants were accepted to attend. The cohort was notably diverse, with 38% self-identifying as underrepresented in the geoscience community and 26% identifying as female. This short course was the first event on EarthScope’s GeoLab platform, a new experimental cloud-based JupyterLab platform hosted by EarthScope. MsPASS (Wang et al., 2022) was the first application to run parallel processing workflows on the GeoLab platform. Participants were exposed to topics such as using MsPASS to process waveform data in the cloud, managing datasets with a document database, and executing data processing workflows in parallel.
@@ 4.2 In-person Events
Our last SCOPED workshop was a five-day hybrid meeting at the University of Washington in Seattle in May 2024 (https://seisscoped.org/workshop-2024/, last accessed 09/10/2024). About 50 participants, including instructors, attended the workshop in person (example of room layout in Fig. 6), along with a varying number of online participants (on average about 50 per day). The training program was led by the research groups of the SCOPED PIs and Dr. Alice-Agnes Gabriel from the University of California San Diego, supported by several NSF- and European-funded projects. Day 1 covered subjects and practicals with an introduction to HPC and Cloud computing and best practices for developing and maintaining open-source software. Day 2 was dedicated to 2D and 3D wave simulations with SPECFEM packages and a tutorial on introduction to full-waveform inversion by SeisFlows and moment tensor inversions with the MTUQ software. Day 3 focused on 3D dynamic rupture and finite source earthquake simulations with SeisSol and 3D wave simulations and computation of 3D adjoint data sensitivity kernels (Tromp et al., 2005; Bozdağ et al., 2011) for full-waveform inversion on a one-chunk mesh with SPECFEM3D_GLOBE (Komatitsch and Tromp, 2002a, b). Day 4 focused on high-precision earthquake catalogs (Wang et al., 2024), where they combined machine learning algorithms (QuakeFlow; Zhu et al. (2023)) with large scale cross-correlation and double-difference methods (HypoDD; Waldhauser and Ellsworth (2000); Waldhauser and Schaff (2008) and demonstrated the use of unsupervised machine learning (SpecUFEx; Holtzman et al. (2018); Sawi et al. (2022)). The participants also had a session on MsPASS (Wang et al., 2022) to learn how to manage big data on HPC and the Cloud. Day 5 addressed ambient noise seismology on the cloud (NoisePy; Jiang and Denolle (2020)), and the trainees were given tutorials on machine learning workflows for seismology on the cloud. All the workshop tutorials were prepared in Jupyter notebooks hosted on GitHub, containerized versions of the open-source SCOPED software were used, and lectures were recorded and uploaded to the SCOPED YouTube channel. All 3D simulations on Day 3 were performed by the trainees on the Frontera system, and the observational seismology tutorials were on AWS.
4.3 Learning Outcomes
@@
Surveys following the SSA and HPS CyberTraining workshops enabled us to evaluate learning outcomes. In the SSA survey, we evaluated the learning outcomes of each module, which were about cloud computing and research-grade applications in ambient noise seismology and machine learning in earthquake catalog building. Eleven survey respondents out of eighty participants noted improved cloud computing skills and a more moderate improvement in disciplinary topics. Overall, the surveyed participants had a positive learning outcome, and improvements in workshop materials and delivery will enhance the impacts of the training.
@@
In the HPS CyberTraining survey, 23 participants responded and expressed positive learning outcomes, with 70% ranking their satisfaction 5/5 and 62.5% indicating that the workshop was a valuable use of their time (rank 5/5). Positive learning outcomes were on Docker and reproducible & open science, frontier seismological topics, and HPC and cloud computing. Several participants expressed verbally or via the survey, that the pace was fast and that instructors should slow down when going through code blocks in notebooks, along with improving participant-led exercises in the notebooks with empty cells.
@@ 5 A Guide to Advanced Computing Workshops
The development of teaching materials requires dedicated effort from both faculty and participants. In-person workshops need a participant-to-assistant ratio of about 15 to 1 for effective debugging. Creating a new email address for event management may add more load. Recruiting participants with similar technical skill levels ensures consistent progress, or additional instruction time can be provided for beginners..
@@ 5.1 Surveys
Evaluation surveys can be useful for quantifying learning outcomes, and crafting them with consideration can benefit professional educators and evaluators. Employing a more standardized approach to the surveys may improve their usefulness. For example, metrics such as “None, Little, Moderate, Quite a bit, Complete” for levels 1 through 5 are similar to the Likert 6-point, “strongly disagree, disagree, somewhat disagree, somewhat agree, agree, strongly agree.” More standard metrics, such as the Likert 6-point, will be incorporated in future surveys to provide a more nuanced measure of fields (Huppenkothen et al., 2018). It is also important to allow only one response per question, as multiple answers can hinder post-event quantitative analysis.
@@ 5.2 Content
Complete virtual meetings can be spread over multiple days and time zones to help attract foreign participants. Virtual meetings can be topical, even narrowly topical, with two sessions: one 2–3 hour session to motivate the science and present a problem for rapid execution of the software and a second 3-hour session for participant-led practice to run the software at scale. This worked well for the SPECFEM and MTUQ workshops.
@@ 5.3.1 Instructor Coordination
Coordination of the workshop instructional team is extremely valuable for the instructors themselves and ultimately improves the flow and delivery of the workshop. We had various degrees of coordination, mostly strong within the PI’s group but less so in the multi-PI workshops. Organizing a workshop, even virtual, requires additional administrative tasks prior to and after the workshop, which were not fully supported for these workshops. The team of instructors should meet frequently and perform at least one full-scale mock event prior to the actual event. We find that the most successful events had a ratio of preparation time over delivery time of 2 or 3, even for workshops that used the same platform and software throughout. It is important to meet frequently with all instructors to build a unified cohort and ensure smooth transitions and delivery during the workshop. Organizers are encouraged to ask non-participating students and colleagues to join mock events and try to follow instructions prior to the workshop in an effort to catch major issues that might otherwise cause undue delays.
@@ 5.3.2 Materials & Platforms
We recommend using Jupyterbooks or Google Docs as shareable platforms to organize the workshop schedule and share training materials, with a clear first page with the schedule that links to the sources (e.g., YouTube recording, Zoom links, GitHub repositories, etc.). Communication platforms like Slack or Teams, or other forms of group communication and direct messaging, allow for rapid, practical communication among instructors and between participants and organizers. Pre-workshop materials are particularly valued and desired by participants. Examples of such materials may be background knowledge, resources, and recorded videos of previous workshops, such as the fundamentals necessary to have acquired by the start of the workshop. Instructors may pre-download data, e.g., pre-processed data and static visualization, to ensure the workshop runs smoothly even with network connectivity of unforeseen technical issues.
@@ 5.3.3 Accounts
The participants’ accounts on the computing resources should be set up days in advance. We provide guidance to automatically create user accounts for AWS on the HPS book. It is important to remind participants that workshop computing resources are temporary. Educational allocations at HPC centers are typically provided when supercomputer center research scientists are involved in the workshop. We recommend assigning the username as the prefix of the participant’s email address, as well as a single generic password, to avoid manual and complicated intervention. Cloud accounts can be created at any time and managed during the workshop. For instance, a cloud manager can re-assign policies, roles, and temporary passwords during the event if needed. Through surveys and emails and possibly “Day 0” virtual help sessions, instructors should ensure that computing setups (accounts, software containers) are working in advance for all virtual participants.
@@ 5.4 First Day
This is the day to onboard participants, ensure that the accounts to HPC and the Cloud are set up and accessible, install ancillary software, and download and test workshop containers to ensure they perform as expected on the participants’ platforms. These tasks can also be done prior to the workshop to free up actual workshop time. Hybrid workshops can be challenging to deliver. They require multiple cameras and microphones for large rooms, attention to remote attendance, and interaction with remote participants. For hybrid events, organizers should ensure that there are sufficient staff/instructors online who can help manage remote participants. Engagement can be improved with frequent polling.
@@ 5.5 During the Event
The feedback on workshops has been positive, especially for focused, single-tool, and single-platform sessions. We found that spending 3 hours per module was reasonable, but 2 hours was too dense and didn’t allow for independent exploration. For virtual workshops, a helpful strategy is to ask "Are you ready to move on?" with the options "Yes, Almost, and No." This helps pace the session and provides instant feedback on participant experience, showing engagement levels and areas needing assistance. Research-grade software often includes advanced workflows tailored to specific use cases, making it challenging to strike a balance between teaching fundamental concepts and realistic scenarios. Some tutorials include additional cells in the notebooks so that participants can test various parameters independently, which was implemented in SPECFEM, MTUQ, and SeisSol. Teaching cloud and HPC computing strategies can be challenging for participants if the content is not relevant to their work. Coordination among instructors with diverse research backgrounds is crucial for effective delivery. One example is to adapt a homegrown software stack based on a specific platform to provision a cloud instance.
@@ 5.6 Post Event
Surveying the participants is a good way to measure learning outcomes. We found that participants will mostly not fill out post-event surveys unless asked at the time of the events, both for virtual and in-person meetings. To improve on evaluation, the exit survey should have similar questions to the incoming survey. Leaving an empty box at the end permits participants to speak freely of things that worked and things to improve.

### v2

@@ 3.4 Open Education
We are compiling a dynamic textbook titled “High-Performance Seismology" (HPS) that the workshop instructors have contributed to HPS (SeisSCOPED, 2024).
@@ 4.1 Virtual Events
In April 2022, we organized a two-day workshop on moment tensor estimation using the open-source MTUQ software (Thurin et al., 2023). The first day featured a 2-hour session introducing key concepts and tutorials. The second day consisted of a 4-hour session that demonstrated how to calculate a library of Green’s functions for a specified 1D layered model using a frequency-wavenumber code (Zhu and Rivera, 2002) and obtain a seismic moment tensor solution. Attendance was strong, with 78 on day 1 and 68 on day 2, indicating sustained interest in the more detailed content. In preparation, software containers for four systems (Windows/PC, Linux, Mac OS Intel, and Mac OS Apple Silicon processors) were developed and tested, resulting in high success rates for participants running the examples.
@@
Building on the success of the previous workshop, we held a three-day SPECFEM users’ workshop in October 2022. Each of the three daily, 4-hour sessions had a specific focus: the forward wavefield (day 1), sensitivity kernels (day 2), and seismic imaging (day 3). Each session included short (20-minute) science lectures, 45-minute tutorials that participants could run locally using pre-downloaded software containers, and wrap-up discussion sessions. Participation ranged from 187 attendees in the day 1 opening seminar to 63 in the day 3 discussion (Figure 5).
This was the first SPECFEM workshop featuring seismic imaging, providing a natural progression from synthetic seismograms (day 1) to sensitivity kernels (day 2) to iterative tomographic inversion using SeisFlows and Pyatoa (day 3) (Modrak et al., 2018; Chow et al., 2020). Crafting a pedagogical but research-grade notebook took a dedicated effort, given the differences in objectives (performance and robustness vs clarity and interactivity). We took some steps to exemplify but downsize large-scale processes in order to retain the same outcome. On the big-data analysis, this meant choosing the duration of the experiment (e.g., 1 day of data) or the spatial extent (e.g., number of stations) to be feasible on 2-4GB worth of RAM, but using the same software tools for the 100 times bigger scale.
@@
In 2023, we held a four-day virtual training workshop in collaboration with the SCEC and several European and NSF-funded projects. Each day focused on a specific theme, starting with an opening day of lectures on open science, reproducibility, software best practices, and an introduction to HPC and Cloud Computing. Subsequent days were divided into subdisciplines and platforms. The workshop attracted over 200 interested participants, with 80 joining on Zoom at the workshop’s start, though attendance varied due to time zone challenges. As the workshop was designed for tool adoptions and relatively fast-paced, participants were exposed to diverse topics in seismology, including earthquake simulations focusing on dynamic rupture (with SeisSol, Käser et al. (2010); Uphoff et al. (2024)) and wave propagation (SPECFEM, Komatitsch et al. (2002)), machine learning phase picking (ELEP, Yuan et al. (2023)), earthquake probabilistic forecasting (pyCSEP, Savran et al. (2022)), and user access to Community Earth Models maintained by SCEC, (e.g., Plesch et al., 2007; Small et al., 2017).
@@
The 2024 MsPASS training short course was hosted in collaboration with EarthScope during the week of July 8, 2024, as part of their 2024 Technical Short Course series. The event featured two hours of lectures and hands-on sessions over three days. Participants could attend the course in real-time or access recordings on YouTube afterward. Daily homework assignments were given, and an optional final project. The application-based enrollment process received 99 valid applications, from which 53 participants were accepted to attend. The cohort was notably diverse, with 38% self-identifying as underrepresented in the geoscience community and 26% identifying as female. This short course was the first event on EarthScope’s GeoLab platform, a new experimental cloud-based Jupyter Lab platform hosted by EarthScope. MsPASS (Wang et al., 2022) was the first application to run parallel processing workflows on the GeoLab platform. Participants were exposed to topics such as using MsPASS to process waveform data in the cloud, managing datasets with a document database, and executing data processing workflows in parallel. 6 participants completed all of the homework assignments, 4 had partial submissions, and 1 completed the optional project.
@@ 4.2 In-person Events
Our most recent SCOPED workshop was a five-day hybrid meeting at the University of Washington in Seattle in May 2024 (https://seisscoped.org/workshop-2024/, last accessed 09/10/2024). About 50 participants, including instructors, attended the workshop in person (example of room layout in Fig. 6), along with a varying number of online participants (on average about 50 per day). The training program was led by the research groups of the SCOPED PIs and Dr. Alice-Agnes Gabriel from the University of California San Diego, supported by several NSF- and European-funded projects. Day 1 covered subjects and practicals with an introduction to HPC and Cloud computing and best practices for developing and maintaining open-source software. Day 2 was dedicated to 2D and 3D wave simulations with SPECFEM packages and a tutorial on introduction to full-waveform inversion by SeisFlows and moment tensor inversions with the MTUQ software. Day 3 focused on 3D dynamic rupture and finite source earthquake simulations with SeisSol and 3D wave simulations and computation of 3D adjoint data sensitivity kernels (Tromp et al., 2005; Bozdağ et al., 2011) for full-waveform inversion on a one-chunk mesh with SPECFEM3D_GLOBE (Komatitsch and Tromp, 2002a, b). Day 4 focused on high-precision earthquake catalogs (Wang et al., 2024), where they combined machine learning algorithms (QuakeFlow; Zhu et al. (2023)) with large scale cross-correlation and double-difference methods (HypoDD; Waldhauser and Ellsworth (2000); Waldhauser and Schaff (2008) and demonstrated the use of unsupervised machine learning (SpecUFEx; Holtzman et al. (2018); Sawi et al. (2022)). The participants also had a session on MsPASS (Wang et al., 2022) to learn how to manage big data on HPC and the Cloud. Day 5 addressed ambient noise seismology on the cloud (NoisePy; Jiang and Denolle (2020)), and the trainees were given tutorials on machine learning workflows for seismology on the cloud. All the workshop tutorials were prepared in Jupyter Notebooks hosted on GitHub, containerized versions of the open-source SCOPED software were used, and lectures were recorded and uploaded to the SCOPED YouTube channel. All 3D simulations on Day 3 were performed by the trainees on the Frontera system, and the observational seismology tutorials were on AWS.
4.3 Learning Objectives and Outcomes
@@
Our main learning objectives were for participants to 1) be able to explain the fundamental principles of high-performance and cloud computing in seismology, 2) apply appropriate computing resources (e.g., HPC clusters, AWS instances) to execute research workflows, 3) compare different computational strategies for seismological research (e.g., traditional local computing vs. cloud/HPC-based approaches), and 4) evaluate their efficiency in handling large-scale seismological data. Surveys following the SSA and HPS CyberTraining workshops enabled us to evaluate some of these learning outcomes. In the SSA survey, we evaluated the learning outcomes of each module, which were about cloud computing and research-grade applications in ambient noise seismology and machine learning in earthquake catalog building. Eleven survey respondents out of eighty participants noted improved cloud computing skills and overall self-reported positive learning outcomes with the workshop. Future improvements in workshop materials and delivery mechanisms will enhance the impacts of the training.
@@
In the HPS CyberTraining survey, 23 participants responded and expressed positive learning outcomes, with 70% ranking their satisfaction 5/5 and 62.5% indicating that the workshop was a valuable use of their time (rank 5/5). Positive learning outcomes were on Docker and reproducible & open science, frontier seismological topics, and HPC and cloud computing. Several participants expressed verbally or via the survey that the pace was fast and that instructors should slow down when going through code blocks in notebooks, along with improving participant-led exercises in the notebooks with empty cells.
@@ 5 A Guide to Advanced Computing Workshops
The development of teaching materials requires dedicated effort from both faculty and participants. In-person workshops need a participant-to-assistant ratio of about 15 to 1 for effective debugging. Recruiting participants with similar technical skill levels ensure consistent progress or additional instruction time can be provided for beginners.
@@ 5.1 Surveys
Evaluation surveys can be helpful in quantifying learning outcomes, and crafting them with consideration can benefit professional educators and evaluators. Employing a more standardized approach to the surveys may improve their usefulness. For example, metrics such as “None, Little, Moderate, Quite a bit, Complete” for levels 1 through 5 are similar to the Likert 6-point, “strongly disagree, disagree, somewhat disagree, somewhat agree, agree, strongly agree.” More standard metrics, such as the Likert 6-point, will be incorporated in future surveys to provide a more nuanced measure of fields (Huppenkothen et al., 2018). Allowing only one response per question is essential, as multiple answers can hinder post-event quantitative analysis.
@@ 5.2 Content
Taken together, these pedagogies may form a module, a self-contained unit that includes 1) a brief lecture (20-30 minutes) introducing key concepts and their relevance to scientific applications, 2) a hands-on tutorial (60-90 minutes) where participants apply the concepts through structured exercises, such as running computations on cloud platforms, setting up HPC environments, or analyzing seismic datasets, and 3) a guided exploration and Q&A (30-60 minutes) to allow participants to troubleshoot their workflows and gain deeper insights. Modules form half-day activities. In-person workshops may combine two modules per day, with 3 hour sessions striking a good balance between depth and independent exploration without too much cognitive fatigue; virtual workshops may spread these modules over multiple days and time zones to help attract foreign participants. This strategy worked particularly well for the SPECFEM and MTUQ workshops.
@@ 5.3.1 Instructor Coordination
Instructor coordination is critical for workshop effectiveness, which was not consistently undertaken in our workshops. We recommend pre-workshop planning meetings to define learning objectives, align instructional materials, and anticipate challenges, mock run-throughs with non-participating students or colleagues to identify unclear instructions and potential bottlenecks in execution, role assignment among instructors, enduring that each focuses on specific tasks such as concept explanation, hands-on support, or software troubleshooting, and diverse instructor background to represent different expertise area (e.g., computational scientists in HPC/Cloud, seismologists form scientific applications, software developers) to allow for a more comprehensive learning experience.
@@ 5.3.2 Materials & Platforms
We used Jupyter Books or Google Docs as shareable, open platforms to organize the workshop schedule and share training materials, with a clear first page with the schedule that links to the sources (e.g., YouTube recording, Zoom links, GitHub repositories, etc.). To-date, our teaching materials are still available and youtube videos still watched. Communication platforms like Slack or Teams, or other forms of group communication and direct messaging, allow for rapid, practical communication among instructors and between participants and organizers. Pre-workshop materials (e.g., pre-requisite tutorials, recorded lectures and videos from previous workshops, and software installation guides) help participants familiarize themselves with foundational concepts beforehand. Instructors may pre-download data, e.g., pre-processed data and static visualization, to ensure the workshop runs smoothly even with unforeseen technical issues such as loss of network connectivity. To accommodate different learning styles, workshops may provide slides and annotated code for visual learners, interactive coding exercises for experiential learners, and may incorporate discussion-based problem-solving for verbal learners.
@@ 5.3.3 Accounts
We found that the workshop ran more smoothly when participants’ accounts on the computing resources was set up days in advance. We provide guidance to automatically create user accounts for AWS on the HPS book. It is important to remind participants that workshop computing resources are temporary. Educational allocations at HPC centers are typically provided when supercomputer center research scientists are involved in the workshop. For our workshop, users chose a simple username, for example, the participants’ email address or its prefix (e.g., <yourID>email.edu), as well as a single generic password to avoid manual and complicated intervention. Cloud accounts can be created at any time and managed during the workshop. For instance, a cloud manager can re-assign policies, roles, and temporary passwords during the event if needed. Through surveys and emails and possibly “Day 0” virtual help sessions, instructors may find it useful to ensure that computing setups (accounts, software containers) are working in advance for all virtual participants.
Participants were made aware that these accounts were temporary and provided with guidelines on how to access these platforms in the future.
@@ 5.4 First Day
This is the day to onboard participants, ensure that the accounts to HPC and the Cloud are set up and accessible, install ancillary software, and download and test workshop containers to ensure they perform as expected on the participants’ platforms. These tasks can also be done prior to the workshop to free up actual workshop time. Hybrid workshops can be challenging to deliver. They require multiple cameras and microphones for large rooms, attention to remote attendance, and interaction with remote participants. For hybrid events, organizers may find it beneficial ensure that there are sufficient staff/instructors online who can help manage remote participants. Engagement can be improved with frequent polling.
@@ 5.5 During the Event
The feedback on workshops has been positive, especially for focused, single-tool, and single-platform sessions. For virtual workshops, a helpful strategy is to ask, "Are you ready to move on?" with the options "Yes, Almost, and No." This helps pace the session and provides instant feedback on participant experience, showing engagement levels and areas needing assistance.
Some of the tutorials, especially those for the core software, included additional cells in the Jupyter Notebooks so that participants could test various parameters independently, which was implemented in SPECFEM, MTUQ, and SeisSol. Other full-stack, research-grade tutorials (e.g., ML-aided earthquake catalog building or ambient noise seismology) included advanced workflows tailored to specific use cases, making it challenging to strike a balance between teaching fundamental concepts and realistic scenarios.
Teaching cloud and HPC computing strategies for research-grade analysis can be challenging for participants if the content is not relevant to their work. In several tutorials, we chose to demonstrate how to adapt a homegrown software stack based on a specific platform to provision a cloud instance. Such an approach allows researchers to upload and deploy their own software stacks on the cloud, ensuring flexibility and independence without imposing a specific platform or software style. While this requires coordination among instructors, it enables each researcher to bring up their preferred tools, reflecting the natural workflow of scientific research.
@@ 5.6 Post Event
Surveying the participants is a good way to measure learning outcomes. We found that most participants will not fill out post-event surveys unless asked at the time of the events, both for virtual and in-person meetings. To improve on the evaluation, the exit survey may benefit from having similar questions to the incoming survey. Leaving an empty box at the end permits participants to speak freely of things that worked and things to improve.